OpenStudy (anonymous):

The base of a triangle is 5/8 feet. The height is 1/3 feet. How much greater is the base of the triangle than the height? A. 7/24 ft B. 1/3 ft C. 4/5 ft D. 23/24 ft

OpenStudy (anonymous):

?? whats the numbers?

OpenStudy (anonymous):

how much feet is the base of the triangle and the height?

OpenStudy (anonymous):

i edited it

OpenStudy (anonymous):

you don't know how to subtract fractions?

OpenStudy (anonymous):

how old are you?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

it's D

OpenStudy (anonymous):

A

OpenStudy (anonymous):

c, d, and a?

OpenStudy (anonymous):

next time you want someone to do the work for you, say in the description.

OpenStudy (anonymous):

c is totally wrong

OpenStudy (anonymous):

dont you subtract the base by the height to find the difference and see how much greater it is

OpenStudy (anonymous):

i want them to help me but they keep giving me the answers

OpenStudy (anonymous):

Sorry my answer is A lol

OpenStudy (anonymous):

5/8 - 1/3 = 15/24 - 8/24 = 7/24

OpenStudy (anonymous):

look up subtracting fractions on google. open one of the links it gives you and enjoy.

OpenStudy (anonymous):

We have subtracted the base and the height to get the answer

OpenStudy (anonymous):

if you want help, it's generally bad policy to swear at people who are here to help you...

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 }-\frac{ 1 }{ 3 }=\frac{ 15 }{ 24 }-\frac{ 8 }{ 24 }=\frac{ 7 }{ 24 }\]

OpenStudy (anonymous):

She did not swear :b

OpenStudy (anonymous):

in pm. twice. and here but then she removed it.

OpenStudy (anonymous):

Oh. You can report if you have any remaining evidence :)

OpenStudy (anonymous):

ignorance is acceptable here. rudeness not so much.

OpenStudy (anonymous):

remember for next time.

OpenStudy (anonymous):

You may tell @hartnn about this

OpenStudy (anonymous):

pm quote: studyislandhelp "report me so what? i'll just make a new account lol"

hartnn (hartnn):

@studyislandhelp please be nice to other users on Open Study. The 1st and the most important point in the code of conduct is to 'Be Nice.' http://openstudy.com/code-of-conduct if that happens again, you may get in trouble. Thanks.

OpenStudy (derpisthewordorisit):

LOOOOOOOOOL!!!!